# Even-parity stability of hairy black holes in $U(1)$ gauge-invariant scalar-vector-tensor theories

### v2

@@ II Background equations in $U(1)$ gauge-invariant SVT theories
From Eq. (15), the current $J_{A}$ of vector field is conserved by virtue of the $U(1)$ gauge symmetry. In addition, if we demand the shift symmetry of the scalar field, the current $J_{\phi}$ will also be conserved. In such a case, this current vanishes at the horizon and hence it should be zero everywhere as we will see in Sec. II. We note that the coupling $\tilde{f}_{3}$ never appears in Eqs. (12)-(15) due to the underlying symmetry of background spacetime.
@@ IV.1 Second-order action and perturbation equations of motion
The coefficients $a_{1}$, $a_{2}$, …, $v_{9}$ are given in Appendix A. In Ref. Kase and Tsujikawa (2023), the even-parity perturbations in Horndeski theories with the interaction between scalar and Maxwell fields characterized by the Lagrangian $G_{2}(\phi,X,F)$ are studied. Compared to that, the presence of the $U(1)$ GI SVT interactions characterized by $f_{3}$, $\tilde{f}_{3}$, $f_{4}$, and $\tilde{f}_{4}$ in Eq. (1) gives rise to the new terms with the coefficients $v_{6}$, $v_{11}$, $v_{12}$, and $v_{13}$, in Eq. (51). For $f_{2}=G_{2}(\phi,X,F)$, $f_{3}=0$, $\tilde{f}_{3}=0$, $f_{4}=0$, and $\tilde{f}_{4}=0$, these coefficients identically vanish and the Lagrangian ${\cal L}_{A}$ coincides with that derived in Ref. Gannouji and Baez (2022); Kase and Tsujikawa (2023). Since the Lagrangian ${\cal L}_{A}$ possesses the similar structure to that in Ref. Kase and Tsujikawa (2023), we can resort to the analogous method in order to identify the dynamical vector DOF in the even-parity sector. Let us introduce an auxiliary field $V(t,r)$ and rewrite Eq. (51) as follows,
@@ IV.2.2 Radial and angular Laplacian stability conditions
In summary, we need to consider five additional conditions (75), (80), (91), (93), and (94), to the stability of odd-parity sector. Nevertheless, it would be still very hard to carry out the stability analysis for the most general case. Instead, we shall process to next section and consider several concrete models.
@@ V Application to the concrete models
respectively, where $\beta_{3}$ and $\beta_{4}$ are arbitrary constants under the constraints led by the odd-parity stability analysis (and later will be further confined by the even-parity ones). The other functions are common to all three models and are chosen as $f_{2}=X+F$ plus ${\tilde{f}}_{3}={\tilde{f}}_{4}=0$. In the following, we shall discuss the stability conditions against odd-parity perturbations, Eqs. (33) and (38), and those against even-parity perturbations, Eqs. (75), (81), (91), (93) and (94), for each model. Notice that, for the convenience of readers, some of the main results are summarized at the end of this section in Table 1. One can move to there directly when the concluding remarks are demanded.
@@ V.1 Stability analysis for the Model 1
where we have defined the dimensionless factor ${\tilde{\beta}}_{3}\equiv M_{\rm pl}\beta_{3}/r_{h}^{2}$ and $r_{h}$ stands for the radii of the event horizon so that we have $f(r_{h})=h(r_{h})=0$. Since Eq. (114) shows that the quantity $\mathcal{K}$ is positive for any non-zero value of ${\tilde{\beta}}_{3}$, the third no-ghost condition will always get satisfied777 According to the (3.14) and (3.15) of Heisenberg and Tsujikawa (2018), we will have $f/h\to 1$ as $r\to r_{h}$. Thus, $\mathcal{K}$ won’t be zero at the event horizon so that it has to stay positive. In addition, notice that, here we are assuming $A_{0}^{\prime}\neq 0$.. We note that the background equations (12)-(15) have been used during simplifications and we shall apply the same operation without additional alerts in the following. Regarding Eqs. (14)-(15), we especially notice that the conditions $J_{\phi}(r)=0$ and $J_{A}(r)=\text{constant}\equiv r_{h}^{2}\kappa$ hold as discussed in the end of Sec. II. Thus, we have
@@ V.2 Stability analysis for the Model 2
The solution for $f$ and $h$ are shown in the panel (a) of Fig. 3 in Appendix C. Since the upper limit of ${\tilde{\beta}_{3}}$ is updated to $\approx 0.441$ compared to Ref. Heisenberg et al. (2018), we adopted it with ${\tilde{\beta}_{4}}=0.1$ and $\mu=0.5$. We can find deviation between Model 2 and that of GR. With the solutions of $f$ and $h$ in hand, together with the previously mentioned algebraic expressions for $A_{0}^{\prime}$ and $\phi^{\prime}$, all the background quantities are known. Inserting them into ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$, we further obtain their numerical values. This part of results are exhibited in the panel (b) of Fig. 3 in which ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$ are plotted as functions of $r_{h}/r$ instead of $r/r_{h}$ to show a larger scope. Fig. 3 shows that, both these two quantities stay positive in the whole space outside the horizon. This result further supports that the first and second angular Laplacian stability conditions for Model 2111111 An ideal treatment is to test the positiveness of ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$ using the Monte Carlo method and try to cover the whole allowed region in $\{\mu,{\tilde{\beta}_{3}},{\tilde{\beta}_{4}},\xi\}$ phase space. Nevertheless, as seen from, e.g., Eqs. (132)-(134), the complexity of our problem has already made the calculations quite time-consuming and it is preventing us from running a more complete analysis. In principle, much more computational resources are needed for executing such an analysis. .
@@ V.3 Stability analysis for the Model 3
$\phi^{\prime}$ $=$ $\frac{2hr\tilde{\beta}_{3}A_{0}^{\prime 2}r_{h}^{2}}{M_{\text{pl}}\left(-6h^{2}\bar{\beta}_{4}A_{0}^{\prime 2}r_{h}^{2}+4h\bar{\beta}_{4}A_{0}^{\prime 2}r_{h}^{2}+fr^{2}\right)}\,,$
@@
$c_{\Omega 3{\rm,even}}^{2}$ $=$ $\frac{{{\tilde{K}}_{33}}}{M_{\text{pl}}^{6}r^{4}\sqrt{fh}\left(2h(2-3h)\bar{\beta}_{4}A_{0}^{\prime 2}r_{h}^{2}+fr^{2}\right)^{2}}$ (128)
×$\times 2L\Big{\{}16h^{5}(3h-2)r^{2}r_{h}^{10}\bar{\beta}_{4}^{2}\tilde{\beta}_{3}^{2}\left(2\bar{\beta}_{4}M_{\text{pl}}^{2}+h\left(2r_{h}^{2}\tilde{\beta}_{3}^{2}-3M_{\text{pl}}^{2}\bar{\beta}_{4}\right)\right)A_{0}^{\prime 10}$
$+4fh^{4}r_{h}^{8}\bar{\beta}_{4}\Big{[}108h^{4}\bar{\beta}_{4}^{2}\left(3M_{\text{pl}}^{2}\bar{\beta}_{4}-2r_{h}^{2}\tilde{\beta}_{3}^{2}\right)M_{\text{pl}}^{4}-72h^{3}\bar{\beta}_{4}^{2}\left(12M_{\text{pl}}^{2}\bar{\beta}_{4}-5r_{h}^{2}\tilde{\beta}_{3}^{2}\right)M_{\text{pl}}^{4}$
$+4\bar{\beta}_{4}\left(8M_{\text{pl}}^{2}\bar{\beta}_{4}\left(2\bar{\beta}_{4}M_{\text{pl}}^{2}+r_{h}^{2}\tilde{\beta}_{3}^{2}\right)-3r^{4}\tilde{\beta}_{3}^{2}\right)M_{\text{pl}}^{2}$
$-8h\left(\left(2r_{h}^{2}\tilde{\beta}_{3}^{4}-3M_{\text{pl}}^{2}\bar{\beta}_{4}\tilde{\beta}_{3}^{2}\right)r^{4}+8M_{\text{pl}}^{4}\bar{\beta}_{4}^{2}\left(6\bar{\beta}_{4}M_{\text{pl}}^{2}+r_{h}^{2}\tilde{\beta}_{3}^{2}\right)\right)$
$+3h^{2}\left(\left(2r_{h}^{2}\tilde{\beta}_{3}^{4}-3M_{\text{pl}}^{2}\bar{\beta}_{4}\tilde{\beta}_{3}^{2}\right)r^{4}+8M_{\text{pl}}^{4}\bar{\beta}_{4}^{2}\left(36M_{\text{pl}}^{2}\bar{\beta}_{4}-5r_{h}^{2}\tilde{\beta}_{3}^{2}\right)\right)\Big{]}A_{0}^{\prime 8}$
$-4f^{2}h^{3}r^{2}r_{h}^{6}\Big{[}6h^{2}\bar{\beta}_{4}^{2}\left(13r_{h}^{2}\tilde{\beta}_{3}^{2}-72M_{\text{pl}}^{2}\bar{\beta}_{4}\right)M_{\text{pl}}^{4}+18h^{3}\bar{\beta}_{4}^{2}\left(12M_{\text{pl}}^{2}\bar{\beta}_{4}-5r_{h}^{2}\tilde{\beta}_{3}^{2}\right)M_{\text{pl}}^{4}$
$+\bar{\beta}_{4}\left(3r^{4}\tilde{\beta}_{3}^{2}-8M_{\text{pl}}^{2}\bar{\beta}_{4}\left(8\bar{\beta}_{4}M_{\text{pl}}^{2}+3r_{h}^{2}\tilde{\beta}_{3}^{2}\right)\right)M_{\text{pl}}^{2}$
$+3h\left(\left(r_{h}^{2}\tilde{\beta}_{3}^{4}-M_{\text{pl}}^{2}\bar{\beta}_{4}\tilde{\beta}_{3}^{2}\right)r^{4}+8M_{\text{pl}}^{4}\bar{\beta}_{4}^{2}\left(12\bar{\beta}_{4}M_{\text{pl}}^{2}+r_{h}^{2}\tilde{\beta}_{3}^{2}\right)\right)\Big{]}A_{0}^{\prime 6}$
$+f^{3}h^{2}r^{4}M_{\text{pl}}^{2}r_{h}^{4}\Big{[}-\tilde{\beta}_{3}^{2}r^{4}-288hM_{\text{pl}}^{4}\bar{\beta}_{4}^{2}+24M_{\text{pl}}^{2}\bar{\beta}_{4}\left(4\bar{\beta}_{4}M_{\text{pl}}^{2}+r_{h}^{2}\tilde{\beta}_{3}^{2}\right)$
$+24h^{2}M_{\text{pl}}^{2}\bar{\beta}_{4}\left(9M_{\text{pl}}^{2}\bar{\beta}_{4}-2r_{h}^{2}\tilde{\beta}_{3}^{2}\right)\Big{]}A_{0}^{\prime 4}$
$+2f^{4}hr^{6}M_{\text{pl}}^{4}r_{h}^{2}\left(8\bar{\beta}_{4}M_{\text{pl}}^{2}+r_{h}^{2}\tilde{\beta}_{3}^{2}+h\left(r_{h}^{2}\tilde{\beta}_{3}^{2}-12M_{\text{pl}}^{2}\bar{\beta}_{4}\right)\right)A_{0}^{\prime 2}+f^{5}r^{8}M_{\text{pl}}^{6}\Big{\}}^{2}$
×$\times\Big{\{}4fhr^{2}A_{0}^{\prime 2}r_{h}^{2}\left[h\left(2\tilde{\beta}_{3}^{2}r_{h}^{2}-3\bar{\beta}_{4}M_{\text{pl}}^{2}\right)+2\bar{\beta}_{4}M_{\text{pl}}^{2}\right]$
$-4h^{2}(3h-2)\bar{\beta}_{4}A_{0}^{\prime 4}r_{h}^{4}\left(h\left(2\tilde{\beta}_{3}^{2}r_{h}^{2}-3\bar{\beta}_{4}M_{\text{pl}}^{2}\right)+2\bar{\beta}_{4}M_{\text{pl}}^{2}\right)+f^{2}r^{4}M_{\text{pl}}^{2}\Big{\}}^{-3}\,.$
@@ Appendix A Coefficients in the second-order action of even-parity perturbations
where ${\cal E}_{11}$ and ${\cal E}_{\phi}$ are the background equations given in Eqs. (12) and (14), respectively, and $\alpha_{i}$ $(i=1,2,4,5,6,7)$ are the coefficients in the second-order action for the odd-parity perturbations defined in Eq. (24).

### v3

@@ II Background equations in $U(1)$ gauge-invariant SVT theories
From Eq. (15), the current $J_{A}$ of vector field is conserved by virtue of the $U(1)$ gauge symmetry. In addition, if we demand the shift symmetry of the scalar field, the current $J_{\phi}$ will also be conserved. In such a case, this current vanishes at the horizon and hence it should be zero everywhere as we will see in Sec. II. Interested readers can check these field equations out in Chao Zhang and Ryotaro Kase . We note that the coupling $\tilde{f}_{3}$ never appears in Eqs. (12)-(15) due to the underlying symmetry of background spacetime.
@@ IV.1 Second-order action and perturbation equations of motion
The coefficients $a_{1}$, $a_{2}$, …, $v_{9}$ are given in Appendix A. In Refs. Kase and Tsujikawa (2023) as well as Chao Zhang and Ryotaro Kase , the even-parity perturbations in Horndeski theories with the interaction between scalar and Maxwell fields characterized by the Lagrangian $G_{2}(\phi,X,F)$ are studied. Compared to that, the presence of the $U(1)$ GI SVT interactions characterized by $f_{3}$, $\tilde{f}_{3}$, $f_{4}$, and $\tilde{f}_{4}$ in Eq. (1) gives rise to the new terms with the coefficients $v_{6}$, $v_{11}$, $v_{12}$, and $v_{13}$, in Eq. (51). For $f_{2}=G_{2}(\phi,X,F)$, $f_{3}=0$, $\tilde{f}_{3}=0$, $f_{4}=0$, and $\tilde{f}_{4}=0$, these coefficients identically vanish and the Lagrangian ${\cal L}_{A}$ coincides with that derived in Ref. Gannouji and Baez (2022); Kase and Tsujikawa (2023). Since the Lagrangian ${\cal L}_{A}$ possesses the similar structure to that in Ref. Kase and Tsujikawa (2023), we can resort to the analogous method in order to identify the dynamical vector DOF in the even-parity sector. Let us introduce an auxiliary field $V(t,r)$ and rewrite Eq. (51) as follows,
@@ IV.2.2 Radial and angular Laplacian stability conditions
In summary, we need to consider five additional conditions (75), (80), (91), (93), and (94) [^7] The relative algebraic expressions of all the nine stability conditions mentioned in this section can be found in Chao Zhang and Ryotaro Kase ., to the stability of odd-parity sector. Nevertheless, it would be still very hard to carry out the stability analysis for the most general case. Instead, we shall process to next section and consider several concrete models.
@@ V Application to the concrete models
[…]
@@ V.1 Stability analysis for the Model 1
where we have defined the dimensionless factor ${\tilde{\beta}}_{3}\equiv M_{\rm pl}\beta_{3}/r_{h}^{2}$ and $r_{h}$ stands for the radii of the event horizon so that we have $f(r_{h})=h(r_{h})=0$. Since Eq. (114) shows that the quantity $\mathcal{K}$ is positive for any non-zero value of ${\tilde{\beta}}_{3}$, the third no-ghost condition will always get satisfied888 According to the (3.14) and (3.15) of Heisenberg and Tsujikawa (2018), we will have $f/h\to 1$ as $r\to r_{h}$. Thus, $\mathcal{K}$ won’t be zero at the event horizon so that it has to stay positive. In addition, notice that, here we are assuming $A_{0}^{\prime}\neq 0$.. We note that the background equations (12)-(15) have been used during simplifications and we shall apply the same operation without additional alerts in the following. Regarding Eqs. (14)-(15), we especially notice that the conditions $J_{\phi}(r)=0$ and $J_{A}(r)=\text{constant}\equiv r_{h}^{2}\kappa$ hold as discussed in the end of Sec. II. Thus, we have
@@ V.2 Stability analysis for the Model 2
The solution for $f$ and $h$ are shown in the panel (a) of Fig. 3 in Appendix C. Since the upper limit of ${\tilde{\beta}_{3}}$ is updated to $\approx 0.441$ compared to Ref. Heisenberg et al. (2018), we adopted it with ${\tilde{\beta}_{4}}=0.1$ and $\mu=0.5$. We can find deviation between Model 2 and that of GR. With the solutions of $f$ and $h$ in hand, together with the previously mentioned algebraic expressions for $A_{0}^{\prime}$ and $\phi^{\prime}$, all the background quantities are known. Inserting them into ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$, we further obtain their numerical values. This part of results are exhibited in the panel (b) of Fig. 3 in which ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$ are plotted as functions of $r_{h}/r$ instead of $r/r_{h}$ to show a larger scope. Fig. 3 shows that, both these two quantities stay positive in the whole space outside the horizon. This result further supports that the first and second angular Laplacian stability conditions for Model 2121212 An ideal treatment is to test the positiveness of ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$ using the Monte Carlo method and try to cover the whole allowed region in $\{\mu,{\tilde{\beta}_{3}},{\tilde{\beta}_{4}},\xi\}$ phase space. Nevertheless, as seen from, e.g., Eqs. (132)-(134), the complexity of our problem has already made the calculations quite time-consuming and it is preventing us from running a more complete analysis. In principle, much more computational resources are needed for executing such an analysis. .
@@ V.3 Stability analysis for the Model 3
$\phi^{\prime}$ $=$ $\frac{2hr\tilde{\beta}_{3}A_{0}^{\prime 2}r_{h}^{2}}{M_{\text{pl}}\left(-6h^{2}\tilde{\beta}_{4}A_{0}^{\prime 2}r_{h}^{2}+4h\tilde{\beta}_{4}A_{0}^{\prime 2}r_{h}^{2}+fr^{2}\right)}\,,$
@@
$c_{\Omega 3{\rm,even}}^{2}$ $=$ $\frac{{{\tilde{K}}_{33}}}{M_{\text{pl}}^{6}r^{4}\sqrt{fh}\left(2h(2-3h)\tilde{\beta}_{4}A_{0}^{\prime 2}r_{h}^{2}+fr^{2}\right)^{2}}$ (128)
×$\times 2L\Big{\{}16h^{5}(3h-2)r^{2}r_{h}^{10}\tilde{\beta}_{4}^{2}\tilde{\beta}_{3}^{2}\left(2\tilde{\beta}_{4}M_{\text{pl}}^{2}+h\left(2r_{h}^{2}\tilde{\beta}_{3}^{2}-3M_{\text{pl}}^{2}\tilde{\beta}_{4}\right)\right)A_{0}^{\prime 10}$
$+4fh^{4}r_{h}^{8}\tilde{\beta}_{4}\Big{[}108h^{4}\tilde{\beta}_{4}^{2}\left(3M_{\text{pl}}^{2}\tilde{\beta}_{4}-2r_{h}^{2}\tilde{\beta}_{3}^{2}\right)M_{\text{pl}}^{4}-72h^{3}\tilde{\beta}_{4}^{2}\left(12M_{\text{pl}}^{2}\tilde{\beta}_{4}-5r_{h}^{2}\tilde{\beta}_{3}^{2}\right)M_{\text{pl}}^{4}$
$+4\tilde{\beta}_{4}\left(8M_{\text{pl}}^{2}\tilde{\beta}_{4}\left(2\tilde{\beta}_{4}M_{\text{pl}}^{2}+r_{h}^{2}\tilde{\beta}_{3}^{2}\right)-3r^{4}\tilde{\beta}_{3}^{2}\right)M_{\text{pl}}^{2}$
$-8h\left(\left(2r_{h}^{2}\tilde{\beta}_{3}^{4}-3M_{\text{pl}}^{2}\tilde{\beta}_{4}\tilde{\beta}_{3}^{2}\right)r^{4}+8M_{\text{pl}}^{4}\tilde{\beta}_{4}^{2}\left(6\tilde{\beta}_{4}M_{\text{pl}}^{2}+r_{h}^{2}\tilde{\beta}_{3}^{2}\right)\right)$
$+3h^{2}\left(\left(2r_{h}^{2}\tilde{\beta}_{3}^{4}-3M_{\text{pl}}^{2}\tilde{\beta}_{4}\tilde{\beta}_{3}^{2}\right)r^{4}+8M_{\text{pl}}^{4}\tilde{\beta}_{4}^{2}\left(36M_{\text{pl}}^{2}\tilde{\beta}_{4}-5r_{h}^{2}\tilde{\beta}_{3}^{2}\right)\right)\Big{]}A_{0}^{\prime 8}$
$-4f^{2}h^{3}r^{2}r_{h}^{6}\Big{[}6h^{2}\tilde{\beta}_{4}^{2}\left(13r_{h}^{2}\tilde{\beta}_{3}^{2}-72M_{\text{pl}}^{2}\tilde{\beta}_{4}\right)M_{\text{pl}}^{4}+18h^{3}\tilde{\beta}_{4}^{2}\left(12M_{\text{pl}}^{2}\tilde{\beta}_{4}-5r_{h}^{2}\tilde{\beta}_{3}^{2}\right)M_{\text{pl}}^{4}$
$+\tilde{\beta}_{4}\left(3r^{4}\tilde{\beta}_{3}^{2}-8M_{\text{pl}}^{2}\tilde{\beta}_{4}\left(8\tilde{\beta}_{4}M_{\text{pl}}^{2}+3r_{h}^{2}\tilde{\beta}_{3}^{2}\right)\right)M_{\text{pl}}^{2}$
$+3h\left(\left(r_{h}^{2}\tilde{\beta}_{3}^{4}-M_{\text{pl}}^{2}\tilde{\beta}_{4}\tilde{\beta}_{3}^{2}\right)r^{4}+8M_{\text{pl}}^{4}\tilde{\beta}_{4}^{2}\left(12\tilde{\beta}_{4}M_{\text{pl}}^{2}+r_{h}^{2}\tilde{\beta}_{3}^{2}\right)\right)\Big{]}A_{0}^{\prime 6}$
$+f^{3}h^{2}r^{4}M_{\text{pl}}^{2}r_{h}^{4}\Big{[}-\tilde{\beta}_{3}^{2}r^{4}-288hM_{\text{pl}}^{4}\tilde{\beta}_{4}^{2}+24M_{\text{pl}}^{2}\tilde{\beta}_{4}\left(4\tilde{\beta}_{4}M_{\text{pl}}^{2}+r_{h}^{2}\tilde{\beta}_{3}^{2}\right)$
$+24h^{2}M_{\text{pl}}^{2}\tilde{\beta}_{4}\left(9M_{\text{pl}}^{2}\tilde{\beta}_{4}-2r_{h}^{2}\tilde{\beta}_{3}^{2}\right)\Big{]}A_{0}^{\prime 4}$
$+2f^{4}hr^{6}M_{\text{pl}}^{4}r_{h}^{2}\left(8\tilde{\beta}_{4}M_{\text{pl}}^{2}+r_{h}^{2}\tilde{\beta}_{3}^{2}+h\left(r_{h}^{2}\tilde{\beta}_{3}^{2}-12M_{\text{pl}}^{2}\tilde{\beta}_{4}\right)\right)A_{0}^{\prime 2}+f^{5}r^{8}M_{\text{pl}}^{6}\Big{\}}^{2}$
×$\times\Big{\{}4fhr^{2}A_{0}^{\prime 2}r_{h}^{2}\left[h\left(2\tilde{\beta}_{3}^{2}r_{h}^{2}-3\tilde{\beta}_{4}M_{\text{pl}}^{2}\right)+2\tilde{\beta}_{4}M_{\text{pl}}^{2}\right]$
$-4h^{2}(3h-2)\tilde{\beta}_{4}A_{0}^{\prime 4}r_{h}^{4}\left(h\left(2\tilde{\beta}_{3}^{2}r_{h}^{2}-3\tilde{\beta}_{4}M_{\text{pl}}^{2}\right)+2\tilde{\beta}_{4}M_{\text{pl}}^{2}\right)+f^{2}r^{4}M_{\text{pl}}^{2}\Big{\}}^{-3}\,.$
@@ Appendix A Coefficients in the second-order action of even-parity perturbations
where ${\cal E}_{11}$ and ${\cal E}_{\phi}$ are the background equations given in Eqs. (12) and (14), respectively, and $\alpha_{i}$ $(i=1,2,4,5,6,7)$ are the coefficients in the second-order action for the odd-parity perturbations defined in Eq. (24). Interested readers may check them out in Chao Zhang and Ryotaro Kase .
@@
Interested readers can check out the mathematical expressions appearing in here with Chao Zhang and Ryotaro Kase .